Suppose the private market for apartments is perfectly competitive with no distortions or market failures. The government is contemplating a project to build 200 apartments. Existing analysis indicates that the current market demand for apartments is given as: Qd =5000 −5P and the market supply is: Qs =5P . Quantity is measured as the number of apartments rented per month at price P. You are part of a team hired by the government to undertake a cost-benefit analysis of the project. Your tasks are to:

(i)Estimate the government’s expected revenues from the project under two options:

(a)the government decides to charge the market price for apartments

(b)the government decides to distribute the apartments to low income families at half the market price.

(ii)Provide an estimate of the gross benefits associated with the government’s project given option (a). Does the decision to pursue option b make a difference to your estimate of gross benefits? Briefly explain your answer.
Answer #1

Qd = 5000- 5P

Qs = 5P

For market equilibrium,

Qd = Qd \Rightarrow 5000 - 5p = 5p \Rightarrowp =500

Government revenue under scheme a = 200* 500 = 100000

Govt. revenue under scheme b = 200* 500/2 = 50000

ii) Social welfare benefits are much higher in option b. Government gets much higher revenue from option a but the social benefit is not that high.